A gas in a close container is heated with 10 Joules of energy causing the lid of th container to rise 2 meters with 3 newton net forcr ehat is the total change in energy of the system?
1
Expert's answer
2018-03-31T09:38:08-0400
The total change in energy of the system is
E=Q+W=Q+Fd
E=10+3∙2=16 J.